OpenStudy (anonymous):

Find the real numbers a b c such that the graph of the function y=ax^2+bx+c contains the points (-2,-4),(1,-1),(3,-19)

OpenStudy (amistre64):

there are a few methods which could work ... which ones are you learning?

OpenStudy (anonymous):

Not sure

OpenStudy (amistre64):

itd be useful if you had some knowledge of the material to work off of ...

OpenStudy (anonymous):

a=34/9 b=31/9 c=-74/9

OpenStudy (anonymous):

thats what i calculated may be wrong tho but do u know how to solve it to start with?

OpenStudy (anonymous):

you will replace the values, get 3 equations then equate one of them like c=..... then get the other 2 equations then use elimination or so

OpenStudy (amistre64):

Hey masterxnonos, We have a policy about answer being given at this site. It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach. Please refer to our code of conduct at http://openstudy.com/code-of-conduct for more information. Thank you.

OpenStudy (amistre64):

and thats not what i get as solutions :)

OpenStudy (anonymous):

I know I have to put them in equations just don't know where to start

OpenStudy (amistre64):

i tend to start with a setup that will use all the points, but zero out things that im not interested in: (-2,-4) (1,-1) (3,-19) when x = the given values, i want certain parts to be zero y = c1 (x-1)(x-3) + c2 (x+2)(x-3) + c3 (x+2)(x-1) each x value will zero out one of the constants allowing us to determine the remaining constant but this is time consuming

OpenStudy (amistre64):

for (-2,-4) -4 = c1 (-2-1)(-2-3) + c2 (-2+2)(x-3) + c3 (-2+2)(x-1) -4 = c1 (-3)(-5) -4/15 = c1 etc.... once the constants are known you can expand it into the quadratic to pull off a b and c

OpenStudy (amistre64):

in the end ...(cough cough)... you can dbl check it with the wolf http://www.wolframalpha.com/input/?i=interpolate+%28-2%2C-4%29%2C%281%2C-1%29%2C%283%2C-19%29

OpenStudy (anonymous):

@amistre64 haha srry man im new didnt know that :P but u know explaining the way is just exhausting >.<

OpenStudy (amistre64):

for this question, i totally agree :)

OpenStudy (anonymous):